puppylover123:

A data set has an odd number of elements, and there are 3 elements above the upper quartile. How many elements does the data set have?

lilaspie7:

idk. wish I could help

XxAngelGirlxX:

13: 13*3/4 = 39/4 = 9.75 … rounded up to 10 So, the upper would be the 11th, the 12th and the 13th ¡3 elements!

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XxAngelGirlxX 13: 13*3/4 = 39/4 = 9.75 … rounded up to 10 So, the upper would be the 11th, the 12th and the 13th ¡3 elements! \(\color{#0cbb34}{\text{End of Quote}}\) I'm just gonna ask how did u solve it without the answer options?

XxAngelGirlxX:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes \(\color{#0cbb34}{\text{Originally Posted by}}\) @XxAngelGirlxX 13: 13*3/4 = 39/4 = 9.75 … rounded up to 10 So, the upper would be the 11th, the 12th and the 13th ¡3 elements! \(\color{#0cbb34}{\text{End of Quote}}\) I'm just gonna ask how did u solve it without the answer options? \(\color{#0cbb34}{\text{End of Quote}}\) I done this question B4 thats why

Hoodmemes:

OOh I get it now can you show me the answer options too?

XxAngelGirlxX:

yea How many elements does the data set have? A. 11 B. 13 C. 15

Hoodmemes:

Thank you

XxAngelGirlxX:

please close this question when done <3 Also your welcome @Hoodmemes